1. You measure 29 textbooks' weights, and find they have a mean weight of 34 ounces. Assume the population standard deviation is 2.8 ounces. Based on this, construct a 95% confidence interval for the true population mean textbook weight.
Give your answers as decimals, to two places
2. If n=28,¯x(x-bar)=36, and s=17, construct a confidence interval at a 90% confidence level. Assume the data came from a normally distributed population.
Give your answers to one decimal place.

Answer by Boreal(15235)

95% half-interval is z*sigma/sqrt(n)=1.96*2.8/sqrt(29)=1.02 books
95% CI is (32.98, 35.02) oz
-
the half-interval is z*17/sqrt(28) where z=1.645, so it is 5.28
using the mean from the first part (not clear)
(28.7, 39.3) oz